Suppose the reserve requirement for all banks is 20 percent and the nation’s banks initially have no excess reserves. Suppose you find US$ 1 million of old currency in your grandmother’s attic and deposit it into her bank checking account. Calculate: (14 %)

The change in reserves in your grandmother’s bank.

The change in required reserves in your grandmother’s bank.
The change in excess reserves in your grandmother’s bank.
The maximum amount by which your grandmother’s bank can expand its loans.
The maximum amount by which the entire banking system can expand its loans.
The potential expansion in the nation’s deposits resulting from the initial deposit in your grandmother’s bank.
The potential expansion in M1.
1
Expert's answer
2020-12-23T07:25:14-0500
  • "1 00 000*0.2=200 000" - The change in reserves in your grandmother’s bank.
  • "1 00 000*0.2=200 000" The change in required reserves in your grandmother’s bank.
  • 0 have no excess reserves
  • "1 000 000-0.2\times1 000 000=800 000" The maximum amount by which your grandmother’s bank can expand its loans.
  • "1 000 000-0.2\times1 000 000=800 000" The maximum amount by which the entire banking system can expand its loan
  • 1 000 000 The potential expansion in the nation’s deposits resulting from the initial deposit in your grandmother’s bank.
  • 1 000 000 - The potential expansion in M1.